OpenStudy (ovidia):

find the x-intercepts of the parabola with vertex(-2,-8)and y-intercept(0,4).

OpenStudy (shansha1):

An equation of the parabola is y = A(x-h)2+k, where (h,k) is the vertex. So, y = A(x+2)2 - 8 Since (0,4) is on the graph, 4 = 4A-8 4A = 12 A = 3 So, y = 3(x+2)2-8 To locate the x-intercepts, set y = 0 and solve for x: 0 = 3(x+2)2-8 3(x+2)2 = 8 x+2 = ±√(8/3) x = -2 ± √(8/3) The x-intercepts are (2+√(8/3),0) and (2-√(8/3),0)